Marie is 4 years younger than Julie. In 3 years the product of their ages will be 410. How old are they?
Marie= x-4
Julie= x
julie= (x+3)
marie= x-4+3 = (x-1)
(x+3)(x-1)= 410


Answer by stanbon(57974)

Marie is 4 years younger than Julie. In 3 years the product of their ages will be 410. How old are they?
Marie= x-4
Julie= x
julie= (x+3)
marie= x-4+3 = (x-1)
(x+3)(x-1)= 410
----
x^2+2x-413 = 0
---
Positive solution:
x = [-2 + sqrt(4-4*-413)]/2
-----
x = [-2 + 40.69]/2
-----
x = 19.35 years (Julie's age now)
x-4 = 15.35 years (Marie's age now)
----
Cheers,
Stan H.
============